OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

OpenStudy (cwrw238):

first expand the brackets -2x + 6 >= 5 - x - 3 now add x to both sides -x + 6 >= 2 what do we do next to isolate -x on left side?

OpenStudy (anonymous):

-2(x-3)>=5-(x+3) Multiply -1 by each term inside the parentheses. -2(x-3)>=5-x-3 Subtract 3 from 5 to get 2. -2(x-3)>=2-x Multiply -2 by each term inside the parentheses. -2x+6>=-x+2 Since -x contains the variable to solve for, move it to the left-hand side of the inequality by adding x to both sides. -2x+6+x>=2 Since -2x and x are like terms, subtract x from -2x to get -x. -x+6>=2

OpenStudy (anonymous):

thank you @cwrw238 and @CJVelasco <3

OpenStudy (anonymous):

yw.

OpenStudy (cwrw238):

welcome